Problem 17-37 (LO. 2, 3)

Amber Industries (a sole proprietorship) sold three § 1231 assets during 2018. Data on these property dispositions are as follows:

Asset Cost Acquired Depreciation Sold For Sold On
Rack $100,000 10/10/14 $100,000 $145,000 10/10/18
Forklift 35,000 10/16/15 23,000 3,000 10/10/18
Bin 87,000 03/12/17 31,000 60,000 10/10/18

a. Determine the amount and the character of the recognized gain or loss from the disposition of each asset. If an amount is zero, enter 0.



Asset

Character of
Gain or Loss
Total
Amount of
Gain or Loss
Amount of
§ 1245
Recapture

Balance of
Gain or Loss
Rack Ordinary income and § 1231 gain
Forklift § 1231 loss
Bin Ordinary income

Amber has $ of ordinary income  due to § 1245 recapture. Additionally, she has $ of net § 1231 gain .

Feedback

b. Assuming that Amber has $5,000 nonrecaptured net § 1231 losses from the five prior years, analyze these transactions and determine the amount (if any) that will be treated as a long-term capital gain.
$
Answer #1

Answer:

a) The amount and the character of the gain or loss recognized from the disposition of each asset should be determined as follows:

Asset Character of Gain or Loss Total Amount of Gain or Loss Amount of §1245 Recapture Balance of Gain or Loss
Rack Ordinary income 145,000 145,000
Forklift §1231 loss (9,000) 9,000
Bin §1231 loss (4,000) 4,000


Amber has $145,000 of ordinary income due to §1245 recapture and $145,000 of  §1231 loss

Workings:

Above figures have been calculated in the following manner:

Asset Character of Gain or Loss Total Amount of Gain or Loss Amount of §1245 Recapture Balance of Gain or Loss
Rack Ordinary income =(145,000+100,000)-100,000 145,000 0
Forklift §1231 loss =(23,000+3,000)-35,000 0 9,000
Bin §1231 loss =(31,000+60,000)-87,000 0 4,000

b) Since Amber has a $145,000 ($9,000+5,000) of  §1231 loss, there is no gain to be treated as capital gain. The $145,000 ,  §1245 depreciation recapture gain will be treated as ordinary income and the $145,000 nett  §1231 will treated as an ordinary loss that will be deducted from the AGI.